Session: Jun 2011 (session 2)
Wavelength: 5 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (Morsiani at Sat May 28 15:58:47 2011) Success:

Torun (K. Boerkowski at Mon May 30 06:43:29 2011) Success:
No known problems.

Yebes (Susana Garcia at Mon May 30 08:45:39 2011) Success:
No known problems.

Effelsberg (Uwe Bach at Mon May 30 10:36:51 2011) Minor failures: FS
LCP lost. Due to an error in the script that produces the procedure files for the DBBC only RCP was recorded. dbbcifa (BBC01-02) and dbbcifb (BBC05-06) both contain RCP data.

Westerbork (Geert Kuper at Mon May 30 17:35:48 2011) Success:
No known problems.

Onsala (Michael Lindqvist at Tue May 31 04:29:30 2011) Minor failures:
Lost first hours due to wind. Started observe around 05:15 UTC, still very windy.